You expect the annual real rate of interest to remain constant for the indefinite future at 2.2%. You expect the upcoming year’s inflation to be 3%. What current one-year spot rate should you observe in the market? Please use the exact formula.

Enter in percent form without the percentage sign. Enter to three decimal places.
Answer #1

Current one-year rate = [(1 + Real Rate) * (1 + Inflation Rate)] - 1

= [(1 + 0.022) * (1 + 0.03)] - 1 = 1.05266 - 1 = 0.05266, or 5.266%